OpenStudy (anonymous):

Find the -coordinate of the point where the graph of the function (ln(sqrtx))/x^2 has a horizontal tangent line. what is the x-coordinate

OpenStudy (anonymous):

\[\ln (\sqrt{x})/x^2\]

OpenStudy (anonymous):

first off \[\frac{\ln(\sqrt{x})}{x^2}=\frac{1}{2}\frac{\ln(x)}{x^2}\] now we need the derivative

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so im not sure where to go from here.

OpenStudy (anonymous):

\[\frac{1-2\ln(x)}{2x^3}\] is the derivative

OpenStudy (anonymous):

set numerator = 0 and solve for x

OpenStudy (anonymous):

wait how did you get 1-2ln(x)/2x^3

OpenStudy (anonymous):

\[2\ln(x)=1\] \[\ln(x)=\frac{1}{2}\] \[x=e^{\frac{1}{2}}=\sqrt{e}\]

OpenStudy (anonymous):

quotient rule

OpenStudy (anonymous):

use \[( \frac{f}{g})'=\frac{gf'-fg'}{g^2}\] with \[f(x)=\ln(x), f'(x)=\frac{1}{x}, g(x)=x^2,g'(x)=2x\]

OpenStudy (anonymous):

using 1/2 as a function and ln(x)/x^2 as another function? and why wouldn't that be product rule

OpenStudy (anonymous):

you have a quotient, not a produce the 1/2 is a constant, so just leave it there

OpenStudy (anonymous):

ok that makes sense i see what your doing

OpenStudy (anonymous):

*product

OpenStudy (anonymous):

so the final answer is e^1/2?

OpenStudy (anonymous):

that is what i got, yes

OpenStudy (anonymous):

ok sounds good to me